Accounting2 Chapters 22, 25 Test, Version 3 Name: There are 30 multiple-choice questions. Each correct answer is worth one of 30 total points. You may write on the test. Please put your answers on your Scantron form and hand it and your test in when you are done. Make sure your name is on your Scantron form and on your test. Use the following information to answer Questions (1) and (2) Lakeside Lumber Company incurs variable costs of $7.00 per board foot in producing its Rough Cut lumber, which it sells for $12.00 per board foot. Alternatively, Lakeside can produce and sell Finished Cut lumber at a price of $18.00 per board foot. Finished Cut variable costs are $10.00 per board foot. If Lakeside decides to produce Finished Cut, per month. Lakesides existing fixed costs are $225,000 per month. the company will have to lease additional equipment at a cost of $65,000 (1) Lakesides Rough Cut break-even unit volume is: (a) 35,000 board feet of lumber (b) 18,750 board feet of lumber (c) 22,500 board feet of lumber (d) 45,000 board feet of lumber (2) Lakesides Rough Cut sales are 70,000 board feet per month. Management estimates that sales volume of Finished Cut lumber would be 50,000 board feet per month. Lakeside should: (pick best answer). (a) continue to produce Rough Cut because the company would be sacrificing sizeable unit sales volume if it made the switch to Finished Cut. b) continue to produce Rough Cut because the change to Finished Cut would increase the companys breakeven point. (c) switch to Finished Cut because contribution margin is increased by $50,000. (d) switch to Finished Cut because income is increased by $50,000. (e) continue to produce Rough Cut because income is decreased by $15,000.
media%2F276%2F27654b9f-884d-4133-bc76-ff
Answer #1

1)Break even Sales volume = Fixed Cost/Contribution Margin per Unit

For Rough Cut:

Break Even Sales Volume = 225000/(12-7)

= 225000/5 = 45000 board feet of lumber

i.e. d


2) Profit from Rough Cut = (70,000-45,000)*5 = 125000

Profit from Finished Cut = (18-10)*50,000 – 225000 – 65000

=110,000

Hence, answer is e – Continue to produce Rough Cut because income is decreased by $15000

3) Relevant cost in this case will be Variable cost because no change in fixed cost due to production

Hence, incremental cost = (18-20)*50000 = 100000 decrease

Hence, d

4) c Prefer to make the component regardless of the required unit volume since Relevant cost per unit is less than purchase cost per unit

5) Contribution per Machine hour of A = (12-5)/3 = 2.333

Contribution per Machine hour of B = (16-6)/5 = 2

Hence, maximum production should be of A

Hence, Answer is c Produce only product A since unlimited demand